OpenStudy (myrandaharris9698):

What are the real roots of x^2+10x+16=0? A. 4 & 4 B. 2 & 8 C. -4 & -4 D. -8 & -2

OpenStudy (anthonyym):

You can graph it using desmos or a graphing calculator. The x-intercepts are the zeros (roots) of the graph.

OpenStudy (myrandaharris9698):

I don't have a graphing calculator

OpenStudy (anonymous):

Factor it. (x+8)(X+2)=0. Set each factor equal to zero and solve for x.

alones (alones):

Okay so factor left side of equation. (x+2)(x+8)=0 and set factors equal to 0 x+2=0 or x+8=0 and for ur answer u should get x=−2 or x=−8